OpenStudy (ray_sihota):

need help. A company finds that when the temperature is 12c it uses 100 liters of heating oil per day. When temprature is 5 c it uses 170 liters of heating oil per day. Find an equation (assumed to be liener) for oil usage in terms of temprature. Find the amount of oil that would be used at -5 c

OpenStudy (jack1):

linear equation, so : y = mx + c x = number of litres y = number of degrees solve for m (gradient) m = (y2-y1) / (x2-x1) m= 5-12 / 170 - 100 m = -7 / 70 m = -1/10 = -0.1 (constant) y = mx + c y = 0.1x + c = ur equation solve for oil if temp (y) = -5

OpenStudy (ray_sihota):

Thank you so much

OpenStudy (jack1):

all good, were u able to solve the 2nd half of the question ok from there...?